# LlamaLens: Specialized Multilingual LLM for Analyzing News and Social Media Content

### v1

@@ 6 Conclusion and Future Work
In this study, we propose a specialized model, LlamaLens, focused on news and social media analysis, designed to assist journalists, fact-checkers, and social media analysts. We curated 52 datasets spanning three prominent languages spoken in the Arabian Peninsula region: Arabic, English, and Hindi. To develop LlamaLens, we created an instruction-following dataset based on these curated datasets and fine-tuned the Llama 3.1-8B-Instruct model. Our experiments show that LlamaLens outperforms the SOTA on 16 datasets, performs comparably on 10 datasets, and underperforms on the rest of the datasets. However, on average, LlamaLens and its quantized versions significantly surpasses the Llama-instruct model. Our findings from error analysis suggests that it is important to inject specialized domain and language knowledge to obtain the desired outcome from Llama-instruct. Our future studies include experimenting with different rank orders and focusing on quantized version of the model to make it usable in a low-resource settings.
@@ A.6 Factuality
For the Factuality Detection task, we utilized five datasets: two in Arabic, two in English, and one in Hindi. The Arabic datasets are “Arafacts” Ali et al. (2021) and “COVID19Factuality” Alam et al. (2021b, a). The English datasets are “News_dataset” Dogo et al. (2020) which combines two datasets (true and fake news) into one. The second English Dataset is “Politifact” Misra (2022b). Finally, the Hindi dataset is “fake-news” Patil et al. (2024). The datasets focus on classifying claims as either true or false to assess their factuality. The SOTA for “COVID19Factuality” by Alam et al. (2021b, a). The SOTA for “News_dataset” was achieved using LSVM as a classifier Ahmed et al. (2017), and for “Politifact” by Rangapur et al. (2023). Moreover, Arafacts SOTA was obtained by  Othman et al. (2024). No SOTA was found for the fake-news in the Hindi dataset.
@@ Appendix B Dataset Sizes After Pre-processing
The Tables 3, 4 and 5 present the sizes of the datasets used in this study, after pre-processing. The datasets are categorized by language (Arabic, English, and Hindi) and we report the distribution of training, development test sets, along with the number of labels for each task.
Task Dataset # Labels # Train # Test # Dev
Attentionworthiness CT22Attentionworthy 9 2,470 1,186 1,071
Checkworthiness CT24_T1 2 22,403 1,031 318
Claim CT22Claim 2 3,513 1,248 339
Cyberbullying ArCyc_CB 2 3,145 900 451
Emotion Emotional-Tone 8 7,024 2,009 1,005
Emotion NewsHeadline 7 939 323 160
Factuality Arafacts 5 4,354 1,245 623
Factuality COVID19Factuality 2 3,513 988 339
Harmful CT22Harmful 2 2,484 1,201 1,076
Hate Speech annotated-hatetweets-4-classes 4 210,526 100,565 90,544
Hate Speech OSACT4SubtaskB 2 4,778 1,827 2,048
News Genre Categorisation ASND 10 74,496 21,942 11,136
News Genre Categorisation SANADAkhbarona 7 62,210 7,824 7,824
News Genre Categorisation SANADAlArabiya 6 56,967 7,123 7,120
News Genre Categorisation SANADAlkhaleej 7 36,391 4,550 4,550
News Genre Categorisation UltimateDataset 10 133,036 38,456 19,269
News Credibility NewsCredibilityDataset 2 8,671 2,730 1,426
@@
Offensive Language ArCyc_OFF 2 3,138 900 450
Offensive Language OSACT4SubtaskA 2 4,780 1,827 2,047
Propaganda ArPro 2 6,002 1,326 672
Sarcasm ArSarcasm-v2 2 8,749 2,996 3,761
Sentiment ar_reviews_100k 3 69,998 20,000 10,000
Sentiment ArSAS 4 13,883 3,976 1,987
Stance Mawqif-Arabic-Stance-main 2 3,162 560 950
Stance stance 3 2,652 379 755
Subjectivity ThatiAR 2 2,446 748 467
@@
Task Dataset # Labels # Train # Test # Dev
Checkworthiness CT24_T1 2 22,403 1,031 318
Claim claim-detection 2 23,224 7,267 5,815
Cyberbullying Cyberbullying 6 32,551 9,473 4,751
Emotion emotion 6 280,551 82,454 41,429
Factuality News_dataset 2 28,147 8,616 4,376
Factuality Politifact 6 14,799 4,230 2,116
News Genre Categorisation CNN_News_Articles_2011-2022 6 32,193 5,682 9,663
News Genre Categorisation News_Category_Dataset 42 145,748 41,740 20,899
News Genre Categorization SemEval23T3-subtask1 3 302 83 130
@@
Offensive Language Offensive_Hateful_Dataset_New 2 42,000 5,252 5,254
@@
Offensive/Hate-Speech hate-offensive-speech 3 48,944 2,799 2,802
Propaganda QProp 2 35,986 10,159 5,125
Sarcasm News-Headlines-Dataset-For-Sarcasm-Detection 2 19,965 5,719 2,858
Sentiment NewsMTSC-dataset 3 7,739 747 320
Subjectivity CT24_T2 2 825 484 219
@@
Task Dataset # Labels # Train # Test # Dev
Cyberbullying MC-Hinglish1.0 7 7,400 1,000 2,119
Factuality fake-news 2 8,393 2,743 1,417
Hate Speech hate-speech-detection 2 3,327 951 476
Hate Speech Hindi-Hostility-Detection-CONSTRAINT-2021 15 5,718 1,651 811
Natural Language Inference Natural Language Inference 2 1,251 447 537
Summarization xlsum 70,754 8,847 8,847
Offensive Speech Offensive Speech Detection 3 2,172 636 318
Sentiment Sentiment Analysis 3 10,039 1,259 1,258
@@ Appendix C Instructions Generation
For each task, each dataset and each language, we use two effective closed models, GPT-4o and Claude-3.5 Sonnet, to generate instructions. These instructions will be used to create the final instruction dataset for LLM fine-tuning. We prompt the models to generate these instructions using the prompt in Table 6. For all generated instructions, we append the following suffix to further instruct the LLM to limit its responses to the labels/summary, to simplify post-processing at inference time: Return only the label without any explanation, justification or additional text. Table 7 shows examples of the generated instructions. Note that we only generate instructions for the user role, while we keep the system role fixed to that presented in Table 7.
@@
User We are creating an English instruction-following dataset for a/an [language] dataset called: [dataset name] covering the task of [task]. The user defined the task as follows: [task definition]. For that task, the labels include: [labels]. Write 10 very diverse and concise English instructions. Return the instructions as strings in a list format as follows [].
Table 6: Prompts used to generate instructions through LLMs.

### v2

@@ 6 Conclusion and Future Work
In this study, we propose a specialized model, LlamaLens, focused on news and social media analysis, designed to assist journalists, fact-checkers, and social media analysts. We curated 52 datasets covering Arabic, English, and Hindi, the key languages of the Arabian Peninsula. Using these, we built an instruction-following dataset and fine-tuned the Llama 3.1-8B-Instruct model for LlamaLens. Our experiments show that LlamaLens outperforms the SOTA on 23 datasets, performs comparably on 8 datasets, and underperforms on the rest of the datasets. However, on average, LlamaLens and its quantized versions significantly surpasses the Llama-instruct model. Our findings from error analysis suggests that it is important to inject specialized domain and language knowledge to obtain the desired outcome. Our future studies include experimenting with different rank orders and focusing on quantized version of the model to make it usable in a low-resource settings.
@@ 7 Limitations
Acknowledgments
The work of M. Hasanain is supported by the NPRP grant 14C-0916-210015 from the Qatar National Research Fund part of Qatar Research Development and Innovation Council (QRDI). The findings achieved herein are solely the responsibility of the authors.
@@ A.6 Factuality
For the Factuality Detection task, we utilized five datasets: two in Arabic, two in English, and one in Hindi. The Arabic datasets are “Arafacts” Sheikh Ali et al. (2021) and “COVID19Factuality” Alam et al. (2021b, a). The English datasets are “News_dataset” Dogo et al. (2020) which combines two datasets (true and fake news) into one. The second English Dataset is “Politifact” Misra (2022b). Finally, the Hindi dataset is “fake-news” Patil et al. (2024). The datasets focus on classifying claims as either true or false to assess their factuality. The SOTA for “COVID19Factuality” by Alam et al. (2021b, a). The SOTA for “News_dataset” was achieved using LSVM as a classifier Ahmed et al. (2017), and for “Politifact” by Rangapur et al. (2023). Moreover, Arafacts SOTA was obtained by  Othman et al. (2024). No SOTA was found for the fake-news in the Hindi dataset.
@@ Appendix B Dataset Sizes After Pre-processing
The Tables 4, 5 and 6 present the sizes of the datasets used in this study, after pre-processing. The datasets are categorized by language (Arabic, English, and Hindi) and we report the distribution of training, development test sets, along with the number of labels for each task.
Task Dataset # Labels # Train # Dev # Test
Attentionworthiness CT22Attentionworthy 9 2,470 1,071 1,186
Checkworthiness CT24_T1 2 22,403 1093 500
Claim CT22Claim 2 3,513 339 1,248
Cyberbullying ArCyc_CB 2 3,145 451 900
Emotion Emotional-Tone 8 7,024 1,005 2,009
Emotion NewsHeadline 7 939 160 323
Factuality Arafacts 5 4,354 623 1,245
Factuality COVID19Factuality 2 3,513 339 988
Harmful CT22Harmful 2 2,484 1,076 1,201
Hate Speech annotated-hatetweets-4-classes 4 210,526 90,544 100,565
Hate Speech OSACT4SubtaskB 2 4,778 2,048 1,827
News Genre Categorization ASND 10 74,496 11,136 21,942
News Genre Categorization SANADAkhbarona 7 62,210 7,824 7,824
News Genre Categorization SANADAlArabiya 6 56,967 7,120 7,123
News Genre Categorization SANADAlkhaleej 7 36,391 4,550 4,550
News Genre Categorization UltimateDataset 10 133,036 19,269 38,456
News Credibility NewsCredibilityDataset 2 8,671 1,426 2,730
@@
Offensive Language ArCyc_OFF 2 3,138 450 900
Offensive Language OSACT4SubtaskA 2 4,780 2,047 1,827
Propaganda ArPro 2 6,002 672 1,326
Sarcasm ArSarcasm-v2 2 8,749 3,761 2,996
Sentiment ar_reviews_100k 3 69,998 10,000 20,000
Sentiment ArSAS 4 13,883 1,987 3,976
Stance Mawqif-Arabic-Stance-main 2 3,162 950 560
Stance stance 3 2,652 755 379
Subjectivity ThatiAR 2 2,446 467 748
@@
Task Dataset # Labels # Train # Dev # Test
Checkworthiness CT24_T1 2 22,403 318 1,031
Claim claim-detection 2 23,224 5,815 7,267
Cyberbullying Cyberbullying 6 32,551 4,751 9,473
Emotion emotion 6 280,551 41,429 82,454
Factuality News_dataset 2 28,147 4,376 8,616
Factuality Politifact 6 14,799 2,116 4,230
News Genre Categorization CNN_News_Articles_2011-2022 6 32,193 9,663 5,682
News Genre Categorization News_Category_Dataset 42 145,748 20,899 41,740
News Genre Categorization SemEval23T3-subtask1 3 302 130 83
@@
Offensive Language Offensive_Hateful_Dataset_New 2 42,000 5,254 5,252
@@
Offensive/Hate-Speech hate-offensive-speech 3 48,944 2,802 2,799
Propaganda QProp 2 35,986 5,125 10,159
Sarcasm News-Headlines-Dataset-For-Sarcasm-Detection 2 19,965 2,858 5,719
Sentiment NewsMTSC-dataset 3 7,739 320 747
Subjectivity clef2024-checkthat-lab 2 825 219 484
@@
Task Dataset # Labels # Train # Dev # Test
Cyberbullying MC-Hinglish1.0 7 7,400 318 1,000
Factuality fake-news 2 8,393 5,815 2,743
Hate Speech hate-speech-detection 2 3,327 4,751 951
Hate Speech Hindi-Hostility-Detection-CONSTRAINT-2021 15 5,718 41,429 1,651
Natural Language Inference Natural Language Inference 2 1,251 4,376 447
Summarization xlsum 70,754 2,116 8,847
Offensive Speech Offensive Speech Detection 3 2,172 9,663 636
Sentiment Sentiment Analysis 3 10,039 20,899 1,259
@@ Appendix C Instructions Generation
For each task, each dataset and each language, we use two effective closed models, GPT-4o and Claude-3.5 Sonnet, to generate instructions. These instructions were used to create the final instruction dataset for LLM fine-tuning. We prompt the models to generate these instructions using the prompt in Table 7. For all generated instructions, we append the following suffix to further instruct the LLM to limit its responses to the labels/summary, to simplify post-processing at inference time: Return only the label without any explanation, justification or additional text. Table 8 shows examples of the generated instructions. Note that we only generate instructions for the user role, while we keep the system role fixed to that presented in Table 8.
@@
User We are creating an [INSTRUCT-LANG] instruction-following dataset for a/an [LANG] dataset called: [DATASET] covering the task of [TASK]. The user defined the task as follows: [TASK DEFINITION]. For that task, the labels include: [LABELS]. Write 10 very diverse and concise English instructions. Return the instructions as strings in a list format as follows [].
Table 7: Prompts used to generate instructions through LLMs. INSTRUCT-LANG refers to the language, which can be Arabic, English, or Hindi. LANG also denotes the language, specifically Arabic, English, or Hindi. TASK refers to the task name. For each task, there is a TASK DEFINITION. LABELS refers to dataset-specific labels.
@@ Appendix E Use of AI assistant
Task Dataset Metric SOTA Base L-Lens (Eng.) L-Lens (Native)
Natural Language Understanding (NLU)
Overall 0.137 0.034 0.129 0.130
News Sum xlsum R-2 0.137 0.034 0.129 0.130
Information Extraction & Text Classification
Overall 0.923 0.660 0.869 0.864
News Cat ASND Ma-F1 0.770 0.587 0.919 0.929
News Cat SANADAkhbarona Acc 0.940 0.784 0.954 0.953
News Cat SANADAlArabiya Acc 0.974 0.893 0.987 0.985
News Cat SANADAlkhaleej Acc 0.986 0.865 0.984 0.982
News Cat UltimateDataset Ma-F1 0.970 0.376 0.865 0.880
News Credibility NewsCredibility Acc 0.899 0.455 0.935 0.933
Subjectivity ThatiAR F1_Pos 0.800 0.562 0.441 0.383
Emotion & Sentiment Analysis
Overall 0.786 0.557 0.724 0.718
Emotion Emotional-Tone W-F1 0.658 0.358 0.705 0.736
Emotion NewsHeadline Acc 1.000 0.406 0.480 0.458
Sarcasm ArSarcasm-v2 F1_Pos 0.584 0.477 0.520 0.542
Sentiment ar_reviews_100k F1_Pos 0.681 0.785 0.779
Sentiment ArSAS Acc 0.920 0.603 0.800 0.804
Stance stance Ma-F1 0.767 0.608 0.926 0.881
Stance Mawqif-Arabic-Stance Ma-F1 0.789 0.764 0.853 0.826
Fast-Checking & Verification
Overall 0.689 0.407 0.663 0.671
Att.worthiness CT22Attentionworthy W-F1 0.412 0.158 0.425 0.454
Checkworthiness CT24_T1 F1_Pos 0.569 0.610 0.502 0.509
Claim CT22Claim Acc 0.703 0.581 0.734 0.756
Factuality Arafacts Mi-F1 0.850 0.210 0.771 0.738
Factuality COVID19Factuality W-F1 0.831 0.492 0.800 0.840
Propaganda ArPro Mi-F1 0.767 0.597 0.747 0.731
Hate Speech & Offensive Content
Overall 0.797 0.603 0.774 0.767
Cyberbullying ArCyc_CB Acc 0.863 0.766 0.870 0.833
Harmfulness CT22Harmful F1_Pos 0.557 0.507 0.523 0.535
Hate Speech annotated-hatetweets-4 W-F1 0.630 0.257 0.526 0.517
Hate Speech OSACT4SubtaskB Mi-F1 0.950 0.819 0.955 0.955
Offensive ArCyc_OFF Ma-F1 0.878 0.489 0.877 0.879
Offensive OSACT4SubtaskA Ma-F1 0.905 0.782 0.896 0.882
Table 9: Arabic Capability-Based Evaluation Results. Base: Llama3.1-Instruct. L-Lens: LlamaLens. Eng.: English. Acc: Accuracy. W-F1: Weighted F1. Mi-F1: Micro-averaged F1. Ma-F1: Macro-averaged F1, R-2: ROUGE-2. Cat:Categorization. News Sum: News Summarization. Att.worthiness: Attentionworthiness.
Task Dataset Metric SOTA Base L-Lens (Eng.) L-Lens (Native)
English Capability-Based Evaluation
Overall 0.152 0.074 0.182 0.181
News Sum xlsum R-2 0.152 0.074 0.182 0.181
Information Extraction & Text Classification
Overall 0.817 0.709 0.669 0.593
News Cat CNN_News_Articles Acc 0.940 0.644 0.970 0.970
News Cat News_Category Ma-F1 0.769 0.970 0.824 0.520
News Genre SemEval23T3-ST1 Mi-F1 0.815 0.687 0.241 0.253
Subjectivity CT24_T2 Ma-F1 0.744 0.535 0.642 0.628
Emotion & Sentiment Analysis
Overall 0.835 0.550 0.830 0.834
Emotion emotion Ma-F1 0.790 0.353 0.803 0.808
Sarcasm News-Headlines Acc 0.897 0.668 0.936 0.947
Sentiment NewsMTSC Ma-F1 0.817 0.628 0.751 0.748
Fact-Checking & Verification
Overall 0.708 0.497 0.811 0.823
Checkworthiness CT24_T1 F1_Pos 0.753 0.404 0.942 0.942
Claim claim-detection Mi-F1 0.545 0.864 0.889
Factuality News_dataset Acc 0.920 0.654 0.999 0.999
Factuality Politifact W-F1 0.490 0.121 0.287 0.311
Propaganda QProp Ma-F1 0.667 0.759 0.963 0.973
Hate Speech & Offensive Content
Overall 0.949 0.529 0.870 0.874
Cyberbullying Cyberbullying Acc 0.907 0.175 0.836 0.855
Offensive Offensive_Hateful Mi-F1 0.692 0.814 0.813
Offensive offensive_language Mi-F1 0.994 0.646 0.899 0.893
Offensive & Hate hate-offensive-speech Acc 0.945 0.602 0.931 0.935
Hindi Capability-Based Evaluation
Overall 0.391 0.356 0.369 0.425
NLI NLI_dataset W-F1 0.646 0.633 0.568 0.679
News Sum xlsum R-2 0.136 0.078 0.171 0.170
Emotion & Sentiment Analysis
Overall 0.697 0.552 0.647 0.654
Sentiment Sentiment Analysis Acc 0.697 0.552 0.647 0.654
Fact-Checking & Verification
Overall 0.759 0.994 0.993
Factuality fake-news Mi-F1 0.759 0.994 0.993
Hate Speech & Offensive Content
Overall 0.703 0.518 0.801 0.802
Hate Speech hate-speech-detection Mi-F1 0.639 0.750 0.963 0.963
Hate Speech Hindi-Hostility W-F1 0.841 0.469 0.753 0.753
Offensive Offensive Speech Mi-F1 0.723 0.621 0.862 0.865
Cyberbullying MC_Hinglish1 Acc 0.609 0.233 0.625 0.627
Table 10: Capability-Based Evaluation Results for English and Hindi. Base: Llama3.1-Instruct. L-Lens: LlamaLens. Eng.: English. Acc: Accuracy. W-F1: Weighted F1. Mi-F1: Micro-averaged F1. Ma-F1: Macro-averaged F1, R-2: ROUGE-2. News Sum: News Summarization. Cat:Categorization. NLI: Natural Language Inference.